OpenStudy (anonymous):

Suppose you prove the identity cot θ = csc θ/sec θ by working only on the right side of the equation. What would the final step look like?

OpenStudy (unklerhaukus):

use \(\csc \theta=\frac 1{\sin\theta}\) and \(\sec\theta=\frac{1}{\cos\theta}\)

mathslover (mathslover):

\[\huge{cot\theta=\frac{csc\theta}{sec\theta}}\] \[\huge{cot\theta=\frac{1}{sin\theta}*cos\theta}\] \[\huge{cot\theta=\frac{cos\theta}{sin\theta}}\]

mathslover (mathslover):

\[\huge{\color{violet}{\textbf{hence proved}}}\]

OpenStudy (unklerhaukus):

the final step would be \[\cot \theta=\frac 1{\tan \theta}\]

OpenStudy (maheshmeghwal9):

yes @UnkleRhaukus is right^_^